OpenStudy (anonymous):

Please help :), how do you find what 8^5/3 ?

ganeshie8 (ganeshie8):

\(\large 8^{\frac{5}{3}}\)

OpenStudy (anonymous):

Thankyou :)

ganeshie8 (ganeshie8):

np, so to find that value, write 8 as 2^3, and see if something cancels out :)

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

\(\large 8^{\frac{5}{3}}\) \(\large (2^3)^{\frac{5}{3}}\)

ganeshie8 (ganeshie8):

look can we cancel something ? :)

OpenStudy (anonymous):

3 and 3

ganeshie8 (ganeshie8):

yup !!

ganeshie8 (ganeshie8):

\(\large 8^{\frac{5}{3}}\) \(\large (2^3)^{\frac{5}{3}}\) \(\large (2^\cancel{3})^{\frac{5}{\cancel{3}}}\) \(\large 2^5\)

OpenStudy (anonymous):

got it thanks :)

ganeshie8 (ganeshie8):

np :)